OpenStudy (anonymous):

Medal and fan !!!!Please help I don't get this its about mathematical induction. use mathematical induction to prove that the statement holds for all positive integers. 2+4+6+...+2n=n^2+n can you display all the steps involved with an explanation of how you get the answer because I really want to understand how to do this. thank you.

OpenStudy (cwrw238):

first show that its true for n = k then its true forn = k + 1 its obviously true for n = 1 so if you can prove the first 2 above then its must be true for n = 1,2, 3 and so on

OpenStudy (anonymous):

how would I do that? like is there an equation I could use ( sums of powers of integers equations?????)

OpenStudy (cwrw238):

you also need to find the formula for the nth term of this series in order to do the above

OpenStudy (cwrw238):

what do you think the nth term is?

OpenStudy (anonymous):

2?

OpenStudy (cwrw238):

lwhat is the relationship between n and nth term n nth term 1 1 2 4 3 6 4 8 can you see the pattern?

OpenStudy (cwrw238):

its 2n right?

OpenStudy (anonymous):

first tern is 2 @cwrw238

OpenStudy (anonymous):

what will I do with the 2n then?

OpenStudy (cwrw238):

ok assume that the sum formula is true for n = k then sum of k terms = k^2 + k so the sum of k+1 terms ( which is sum of k terms + (k + 1) th term ) = k^2 + k + 2(k+1) agreed?

OpenStudy (cwrw238):

In my first post i should have said ASSUME true for k not show true for k

OpenStudy (cwrw238):

do you follow the argument so far?

OpenStudy (cwrw238):

nth term = 2n so (k+1)th term = 2(k +1)

OpenStudy (anonymous):

somewhat

OpenStudy (cwrw238):

ok so so far we have worked out that sum of (k+1) terms = k^2 + k + 2(k + 1) k^2 + k + 2(k + 1) = k^2 + k + 2k + 2 =k^2 + 2k + 1 + k + 1 = (k + 1)^2 + k + 1

OpenStudy (cwrw238):

- notice that this sum formula is the same as the formula for the of k terms except that k is replaced by k+1 ok?

OpenStudy (cwrw238):

* same as sum for k terms except that k is replaced by (k +1)

OpenStudy (cwrw238):

ah well!! - he seems to have given up!

OpenStudy (cwrw238):

good you are back do you follow this ok?

OpenStudy (anonymous):

wait who? im lost. so k is replaced by k+1 what happens next to the equation? oh its a she.

OpenStudy (cwrw238):

yep im a she lol

OpenStudy (anonymous):

so am I lol im terrible at math lol

OpenStudy (cwrw238):

if you compare the 2 formula sum of k terms = k^2 + k sum of k+1 terms = (k+1)^2+ k + 1 - from this we can say that if the sum of k formula is true then the sum of (k+1) formula is true right?

OpenStudy (anonymous):

LOL no one is terrible at math its this thought which makes them think they are terrible keep practicing no one is as bad as they think they are.

OpenStudy (cwrw238):

i agree

OpenStudy (anonymous):

yeah k+1 is true. when it comes to calculus I am alright I have to admit but at times I get dumbfounded with these problems because I guess what they ask confuses me

OpenStudy (anonymous):

thats the algorithm for proving induction problems you should probably learn it.

OpenStudy (anonymous):

ik that in order to prove the induction I need to make sure pi is true and the truth of p implies the truth of pk+1 for every positive k , then pn has to be true for all positive integers. I get that to a degree when it comes the expression itself I just get confused with the steps involved.

OpenStudy (cwrw238):

yea Proof by Induction can be confusing at first basically for a sequence it consists of 1 Assume that the sum formula is true for n = k 2 Work out what the nth term is. Then put n = k+1 and add this to the formula for sum formula for n = k 3. Then manipulate this algebraically to try and get the same formula as k except that k+1 replaces k. This proves that if the sum formula is true for k then its true for k+1. 4. Now show that sum formula is true for n = 1. 5. This concludes the proof because its true for k+1 if sum of k is correct and its true for k=1 then its must be true for k=2 and k = 3 and so one. So true for all n.

OpenStudy (anonymous):

since we found k+1 to be true there is no more to do in the equation then?

OpenStudy (cwrw238):

- yes - you have to contrive to get the expression in same form as k expression can be tricky

OpenStudy (cwrw238):

except to show that its true for k =1 which it obviously is in this case. Sum of 2 = 2!

OpenStudy (anonymous):

so sum of 3 and 4 would be 3, 4?

OpenStudy (cwrw238):

no want i meant was sum of 1 term = 2 as 2 is first term sum of 2 terms = 2+4

OpenStudy (anonymous):

oh...hehe.

OpenStudy (cwrw238):

lol

OpenStudy (cwrw238):

test the formula if you like sum of 2 terms = 2^2 + 2 = 6 2 +4 = 6

OpenStudy (anonymous):

okay so the next thing to do would be to add 6+___= to get k=3?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

but you would not do this if you were asked to find sum of 90 terms for instance - it would take a lot of adding you would replace n by 90 in the formula sum of 90 terms = 90^2 + 90 = 8190

OpenStudy (anonymous):

ok so it would be 6+ 8 or 6+12?

OpenStudy (cwrw238):

you mean the sum of 3 terms?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

sorry 2 + 4 +6 = 12

OpenStudy (anonymous):

k=4 would then be 2+4+6+12=?

OpenStudy (cwrw238):

the important thing is for you to try and understand the method of induction and practice some problems

OpenStudy (cwrw238):

no the next term after 6 is 8 - they are going up in 2's

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for the incredible help!

OpenStudy (cwrw238):

yw